OpenStudy (kendricklamar2014):

Solve this system using elimination. Check your answer. 6x + 4y = 42 and -3x + 3y = -6

OpenStudy (anonymous):

We need to eliminate the x or y. How would you do this when combining the two equations?

OpenStudy (anonymous):

What happens if you multiply the second equation by 2 then combine with the first equation?

OpenStudy (phi):

it's often nice to first simplify the equations if you can. if you can divide each of the numbers (by 2 in the first equation, for example), do that first. 6x+4y=42 divide *every term, both sides* by 2 3x + 2y = 21 we could simplify the second equation -3x +3y=-6 but notice what happens if we add the two equations.

OpenStudy (phi):

can you write it down ?

OpenStudy (kendricklamar2014):

Like this: 3x + 2y = 21 + -3x + 3y = -6

OpenStudy (phi):

ok, but you add left side to left side and right side to right side

OpenStudy (kendricklamar2014):

Im confused @phi

OpenStudy (phi):

3x + 2y = 21 -3x + 3y = -6 ------------- <--- sum goes here

OpenStudy (kendricklamar2014):

3x + -3x = 0 2y + 3y = 5y 21 + -6 = 15

OpenStudy (phi):

ok, but put the answer below the lines, and write it as a new equation

OpenStudy (kendricklamar2014):

Like this: 0 + 5y = 15

OpenStudy (phi):

yes

OpenStudy (phi):

you can simplify it because adding 0 can be ignored

OpenStudy (phi):

there is more to do so far you have 5y= 15 (I assume you know 0+5y simplifies to 5y) now divide both sides by 5

OpenStudy (kendricklamar2014):

5y / 5 = 15/5 y= 3

OpenStudy (phi):

next, pick one of the equations you started with. 6x + 4y = 42 replace y with 3. solve for x. can you do that?

OpenStudy (kendricklamar2014):

so like this: 6x + 4*3 = 42 right

OpenStudy (phi):

yes

OpenStudy (kendricklamar2014):

So, then the answer would come out to be X=5 correct

OpenStudy (phi):

you can test x=5, y=3 in the *second* equation. if they work, you found the x and y that "solves" the system

OpenStudy (phi):

-3x + 3y = -6 test with x=5 and y=3 that means replace x with 5, and y with 3 and simplify if you get -6=-6 it works

OpenStudy (kendricklamar2014):

So 6*5 + 4*3 = 42 42=42

OpenStudy (kendricklamar2014):

So its true

OpenStudy (phi):

you have to check that they work for both equations. but, so far, so good

OpenStudy (kendricklamar2014):

So for the other equation: -3x + 3y = -6 we replace x with 5 and y with 3 right

OpenStudy (phi):

yes

OpenStudy (kendricklamar2014):

-3*5 + 3*3 = -6 -6 = -6 True

OpenStudy (phi):

that shows that x=5, y=3 or if we write it as an (x,y) pair, the point (5,3) is the answer

OpenStudy (kendricklamar2014):

Yes, Thanks

OpenStudy (phi):

if you have time, look over how we solved this. it should make some sense.

OpenStudy (kendricklamar2014):

Yeah, im checking how we solved this to take notes

OpenStudy (phi):

if you have time, seee https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-systems-topic/cc-8th-systems-elimination/v/solving-systems-by-elimination and the other videos in this section.

OpenStudy (kendricklamar2014):

Ok, thanks